# On the Analysis of Misspecified Variational Inequalities with Nonlinear Constraints

### v1

@@ 4 Convergence Analysis
Suppose Assumption 2.1 holds, and define $D_{X}\triangleq\sup_{x\in X}\|x-x^{*}\|<\infty$ and $D_{\Theta}\triangleq\sup_{\theta\in\Theta}\|\theta-\theta^{*}\|<\infty$. Consider Algorithm 1 with constant penalty $\rho_{k}\equiv\rho>0$ and primal step-size $\gamma_{k}\equiv\gamma\in\bigl(0,(L^{\Phi}+2L_{F,x}+L_{F,\theta})^{-1}\bigr)$. Let $\alpha_{k}=\tfrac{c}{(k+1)^{1+\delta}}$ for any $c\in(0,\frac{1}{2})$ and $\delta>0$, $\rho\leq\frac{1}{L^{\Phi}_{\lambda\theta}}$, and the learning error satisfying $\sum_{k=0}^{\infty}\alpha_{k}^{-1}\|\theta_{k}-\theta^{*}\|^{2}<\infty$. Then the dual sequence $\{\lambda_{k}\}$ is bounded and convergent; in particular $\|\lambda_{k}-\lambda^{*}\|$ converges to a finite limit.
@@
See Appendix C for the proof. ∎
Lemma 4.3 establishes bounded dual iterates under compactness and summable learning errors when we use a constant primal step-size and a penalty parameter. We adopt this schedule henceforth, which yields a uniform bound $D_{\Lambda}\triangleq\sup_{k\geq 0}\|\lambda_{k}\|<+\infty$. With this standing in place, we now state the main convergence result for the ergodic primal averages.
@@
Suppose Assumption 2.1 holds, and Algorithm 1 is executed under the conditions of Lemma 4.3. Then, for the ergodic average $\bar{x}_{K}=\tfrac{1}{K}\sum_{k=1}^{K}x_{k}$, the following hold:
@@
$\text{where}\quad C^{\prime}_{T}$ $\triangleq\Big(\tfrac{1}{2\gamma}+\tfrac{3}{2}L_{F,x}\Big)D_{X}^{2}+2L_{F,\theta}D_{X}D_{\Theta},$
$C^{\prime}_{R}(x,0)$ $=\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+\rho L_{f,\theta}JD_{f}(\theta^{*})\Big)\ \sum_{k}\|\theta_{k}-\theta^{*}\|$
$\quad+2L_{F,\theta}\sum_{k}\|\theta_{k}-\theta^{*}\|^{2},$
$C_{\mathrm{feas}}$ $=C^{\prime}_{T}+\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+L^{\Phi}_{\lambda\theta}(\sqrt{J}+2D_{\Lambda})+\rho L_{f,\theta}JD_{f}(\theta^{*})\Big)\ \sum_{k}\|\theta_{k}-\theta^{*}\|$
$\quad+2L_{F,\theta}\sum_{k}\|\theta_{k}-\theta^{*}\|^{2}+\frac{(\sqrt{J}+2D_{\Lambda})^{2}}{2\rho}.$
@@
Based on Assumption 2.1-(ii), the secondary VI problem (2) has a Lipschitz continuous and strongly monotone operator which implies that the sequence $\{\theta_{k}\}_{k\geq 0}$ generated by the projected gradient step in line 6 of Algorithm 1 converges to the optimal unique parameter $\theta^{*}$ at a linear convergence rate (e.g. see [7, Proposition 26.16]); hence, the error accumulations $\sum_{k}\|\theta_{k}-\theta^{*}\|$ in the upper bounds of the result of Theorem 4.4 are finite with a uniform bound. Therefore, the result of Theorem 4.4 implies that achieving $\epsilon$-infeasibility and $\epsilon$-relaxed gap requires at most running $\mathcal{O}(1/\epsilon)$ iterations of Algorithm 1. We finally remark that the update of misspecified parameter $\theta$ in the algorithm can be replaced with other methods, such as extra-gradient method, as long as $\theta_{k}\to\theta^{*}$ at a linear rate.
@@ 5 Numerical Experiment
In this experiment, for each player $i\in\{1,\ldots,N\}$ we let the capacity of each firm be $\text{Cap}_{i}=5$ and the quadratic cost function $c_{i}(x_{i})=\tfrac{1}{2}r_{i}x_{i}^{2}+g_{i}x_{i}$ where $r_{i},g_{i}$ are randomly generated from uniform distributions $U[1,10]$ and $U[5,20]$, respectively. Moreover, $p(X;b^{*})=a^{*}-b^{*}X$ where $X=\sum_{i=1}^{N}x_{i}$, we let $a^{*}=100$, and $b^{*}$ is the solution of the least-square minimization (3) with synthetic observations generated from a true model $X_{t}\sim U[2,20]$, $p_{\mathrm{obs},t}=a^{*}-b^{*}X_{t}$, for $t=1,\dots,T=300$. All the firms should also satisfy the coupling constraint of the form $f(\mathbf{x},b^{*})\triangleq p(X;b^{*})-\delta\leq 0$ for $\delta=0.95a^{*}$.
We implemented our proposed ALM-Misspecified VI method and compare it with the misspecified Tikhonov regularized VI [9, 3] and Extragradient (EG) method by [3]. For all methods, the misspecified parameter is updated using one projected gradient step at each iteration. To handle the coupling constraint for the other two methods, we employ the Lagrangian reformulation for each player and construct the corresponding VI operator by concatenating all the players’ gradient vectors of the Lagrangian function with respect to primal and dual variables. We consider three problem instances with $(N,D)\in\{(50,5),(50,10),(100,10)\}$ and run each method for a sufficiently large number of iterations. Figure 1, shows the performance of the methods in terms of the VI gap and infeasibility metrics. Our proposed method (ALM-Misspecified VI) exhibits a superior performance across all instances, while the Lagrangian-Tikhonov and EG-Lagrangian baselines converge more slowly and deteriorate more noticeably as $(N,D)$ grow.
@@
Let $j$ be an index in $\{1,\dots,J\}$. For any $j^{th}$ coordinate using algorithm 1 (line 5), the dual update is $\lambda_{k+1}^{(j)}=\Bigl[\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})\Bigr]_{+},$ where $\cdot$$_{+}$ denotes the projection onto $\mathbb{R}_{+}$. We now consider two cases, when $f_{j}^{\prime}s$ are active, we have $\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})>0.$ Then the projection acts as the identity, so we have
@@
Thus, we conclude that in the active case $\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}=\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}.$ Again, if $f_{j}(x_{k+1},\theta_{k})^{\prime}s$ are inactive, then we have $\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})<0\quad\implies f_{j}(x_{k+1},\theta_{k})<-\frac{\lambda_{k}^{(j)}}{\rho_{k}}.$ Thus, we have $\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}=-\lambda_{k}^{(j)}.$ Therefore, in inactive case also we have $\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}=\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}.$ Since the above equality holds for both active and inactive cases, we can write in full vector notation that $\lambda_{k+1}-\lambda_{k}=\rho_{k}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k}).$ Now by adding and subtracting $\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$, we get
@@
where in the last equality, we define $e_{k}=\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)-\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)$. Let any $\lambda\in\mathbb{R}_{+}^{J}$, then using the three points inequality and dividing both sides by $\rho_{k}$ we get
@@
$\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}$ $=\frac{1}{2\rho_{k}}\|\lambda_{k}-\lambda\|^{2}-\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda\|^{2}+(\lambda_{k+1}-\lambda)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)$
$\quad+(\lambda_{k+1}-\lambda)^{\top}e_{k}$
@@
Squaring both sides and applying the inequality $2ab\leq 2|a||b|$, we obtain
$\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}$ $\geq\rho_{k}^{2}\left[\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-2|f_{j}(x_{k+1},\theta^{*})||\left(f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})\right)|\right].$
@@
$\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}\geq\rho_{k}^{2}\left[\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-2|f_{j}(x_{k+1},\theta^{*})|L_{f,\theta}\|\theta_{k}-\theta^{*}\|\right].$
@@
$\frac{1}{2\rho_{k}}\sum_{j\in J_{k}^{+}}\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}\geq\frac{\rho_{k}}{2}\sum_{j\in J_{k}^{+}}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$
If $j\in J_{k}^{-}$, then $\lambda^{(j)}_{k+1}=0\implies\left(\lambda^{(j)}_{k+1}-\lambda^{(j)}_{k}\right)^{2}=\left(\lambda^{(j)}_{k}\right)^{2}$. Multiplying both side by $\frac{1}{2\rho_{k}}$, and combining both the above inequalities, we get
$\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}\geq\tfrac{\rho_{k}}{2}\sum_{j\in J_{k}^{+}}f_{j}(x_{k+1},\theta^{*})^{2}+\tfrac{1}{2\rho_{k}}\sum_{j\in J_{k}^{-}}(\lambda_{k}^{(j)})^{2}-\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$
@@
$-\left(\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|\right)$
$\qquad\leq-\left[\tfrac{\rho_{k}}{2}\sum_{j\in J_{k}^{+}}f_{j}(x_{k+1},\theta^{*})^{2}+\tfrac{1}{2\rho_{k}}\sum_{j\in J_{k}^{-}}(\lambda_{k}^{(j)})^{2}\right].$ (11)
Also, from the definition we have
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)=\sum_{j\in J_{k}^{+}}\Bigl(\frac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta^{*})\Bigr)-\sum_{j\in J_{k}^{-}}\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}.$ (12)
Now using the facts from (6) and (12), we can expand as follows. First, by substituting the definitions of $\Phi_{\rho_{k}}$, we obtain
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
@@
$\qquad-\sum_{j\in J_{k}^{-}}\bigl(\lambda^{(j)}_{k+1}-\lambda^{(j)}\bigr)\Bigl(-\tfrac{\lambda^{(j)}_{k}}{\rho_{k}}\Bigr)-\sum_{j\in J_{k}^{+}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\sum_{j\in J_{k}^{-}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*}).$
@@
$\quad-\sum_{j\in J_{k}^{+}}\left(\lambda^{(j)}_{k}+\rho_{k}f_{j}(x_{k+1},\theta^{*})-\lambda^{(j)}\right)f_{j}(x_{k+1},\theta^{*})-\sum_{j\in J_{k}^{-}}\left(-\lambda^{(j)}\right)\left(-\frac{\lambda^{(j)}_{k}}{\rho_{k}}\right)$
$\quad-\sum_{j\in J^{+}_{k}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*}).$
@@
$\quad-\left(\lambda^{(j)}_{k}+\rho_{k}f_{j}(x_{k+1},\theta^{*})-\lambda^{(j)}\right)f_{j}(x_{k+1},\theta^{*})\Bigr]-\sum_{j\in J_{k}^{-}}\left(\tfrac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}+\lambda^{(j)}\Bigl(f_{j}(x_{k+1},\theta^{*})+\tfrac{\lambda_{k}^{(j)}}{\rho_{k}}\Bigr)\right).$
@@
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad=-\sum_{j\in J_{k}^{+}}\tfrac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}-\sum_{j\in J_{k}^{-}}\left(\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}+\lambda^{(j)}\Bigl(f_{j}(x_{k+1},\theta^{*})+\frac{\lambda_{k}^{(j)}}{\rho_{k}}\Bigr)\right).$ (13)

### v3

@@ 4 Convergence Analysis
Suppose Assumption 2.1 holds, and define $D_{X}\triangleq\sup_{x\in X}\|x-x^{*}\|<\infty$ and $D_{\Theta}\triangleq\sup_{\theta\in\Theta}\|\theta-\theta^{*}\|<\infty$. Consider Algorithm 1 with constant penalty $\rho_{k}\equiv\rho>0$ and primal step-size $\gamma_{k}\equiv\gamma\in\bigl(0,(\rho C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}+C_{2}D_{\Lambda})^{-1}\bigr)$ for some $D_{\Lambda}>0$ where $C_{1},C_{2}$ are defined in Lemma 4.2. Assume that the learning error satisfies $\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|<\infty$ and $\sum_{k=0}^{\infty}(k+1)^{2}\|\theta_{k}-\theta^{*}\|^{2}<\infty,$ and $\rho_{k}=\rho=\frac{1}{L_{\lambda\theta}^{\Phi}}$ for all $k\geq 0$. Then the dual sequence $\{\lambda_{k}\}$ is bounded.
@@
Proof See Appendix D for the proof. ∎
Lemma 4.4 establishes bounded dual iterates under compactness and summable learning errors when we use a constant primal step-size and a penalty parameter. We adopt this schedule henceforth, which yields a uniform bound $D_{\Lambda}\triangleq\sup_{k\geq 0}\|\lambda_{k}-\lambda^{*}\|<+\infty$. With this standing in place, we now state the main convergence result for the ergodic primal averages.
@@
Suppose Assumption 2.1 holds, and Algorithm 1 is executed under the conditions of Lemma $4.$ Then, for the ergodic average $\bar{x}_{K}=\tfrac{1}{K}\sum_{k=1}^{K}x_{k}$, the following hold:
@@
where $C^{\prime}_{T}\triangleq\tfrac{1}{2\gamma}D_{X}^{2}+2L_{F,\theta}D_{X}D_{\Theta}$, $C^{\prime}_{R}(x,0)=\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}$, and $C_{\mathrm{feas}}=C^{\prime}_{T}+\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}+\frac{J+\|\lambda_{0}-\lambda^{*}\|^{2}}{\rho}$.
@@
Based on Assumption 2.1-(ii), the secondary VI problem (2) has a Lipschitz continuous and strongly monotone operator which implies that the sequence $\{\theta_{k}\}_{k\geq 0}$ generated by the projected gradient step in line 7 of Algorithm 1 converges to the optimal unique parameter $\theta^{*}$ at a linear convergence rate (e.g. see [7, Proposition 26.16]); hence, the learning error accumulations $\sum_{k}\|\theta_{k}-\theta^{*}\|$ and $\sum_{k}(k+2)^{2}\|\theta_{k}-\theta^{*}\|^{2}$ in the upper bounds of the result of Theorem 4.5 are finite with a uniform bound. Therefore, the result of Theorem 4.5 implies that achieving $\epsilon$-infeasibility and $\epsilon$-relaxed gap requires at most running $\mathcal{O}(1/\epsilon)$ iterations of Algorithm 1. We finally remark that the update of misspecified parameter $\theta$ in the algorithm can be replaced with other methods, such as extra-gradient method, as long as $\theta_{k}\to\theta^{*}$ at a linear rate.
@@ 5 Numerical Experiment
We consider a market with $N$ firms, where each firm $i\in\{1,\dots,N\}$ produces $D$ homogeneous products so the total decision variable dimension is $n=ND$. For each product $d=1,\dots,D$ of firm $i$, the capacity is $\text{Cap}_{i}=5$ and the quadratic cost function $c_{i}(x_{i})=\tfrac{1}{2}r_{i}x_{i}^{2}+g_{i}x_{i}$ where $r_{i},g_{i}$ are randomly generated from uniform distributions $U[1,10]$ and $U[5,20]$, respectively. Moreover, $p(X;b^{*})=a^{*}-b^{*}X$ where $X=\sum_{i=1}^{N}x_{i}$, we let $a^{*}=100$, and $b^{*}$ is the solution of the least-square minimization (3) with synthetic observations generated from a true model $X_{t}\sim U[2,20]$, $p_{\mathrm{obs},t}=a^{*}-b^{*}X_{t}$, for $t=1,\dots,T=300$. All the firms should also satisfy the coupling constraint of the form $f(\mathbf{x},b^{*})\triangleq p(X;b^{*})-\delta\leq 0$ for some $\delta>0$.
We implemented our proposed ALM-Misspecified VI method and compare it with the misspecified Tikhonov regularized VI [9, 3] and Extragradient (EG) method by [3]. For all methods, the misspecified parameter is updated using one projected gradient step at each iteration. To handle the coupling constraint for the other two methods, we employ the Lagrangian reformulation for each player and construct the corresponding VI operator by concatenating all the players’ gradient vectors of the Lagrangian function with respect to primal and dual variables. We consider three problem instances with $(N,D)\in\{(50,5),(50,10),(100,10)\}$ and run each method for a sufficiently large number of iterations with constant step-sizes chosen according to the theoretical findings. Figure 1, shows the performance of the methods in terms of the VI gap $\widetilde{\mathrm{Gap}}(\bar{x}_{K},b^{*})$ and infeasibility $\mathbf{1}^{\top}[f(\bar{x}_{K},b^{*})]_{+}$. Our proposed method (ALM-Misspecified VI) exhibits a superior performance across all instances, while the Lagrangian-Tikhonov and EG-Lagrangian baselines converge more slowly and deteriorate more noticeably as $(N,D)$ grow.
@@
Proof Let $j$ be an index in $\{1,\dots,J\}$. For any $j^{th}$ coordinate using Algorithm 1 (line 6), the dual update is $\lambda_{k+1}^{(j)}=\Bigl[\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})\Bigr]_{+},$ where $\cdot$$_{+}$ denotes the projection onto $\mathbb{R}_{+}$. We now consider two cases, when $f_{j}^{\prime}s$ are active, we have $\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})>0.$ Then the projection acts as the identity, so we have
@@
Thus, we conclude that in the active case $\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}=\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}.$ Again, if $f_{j}(x_{k+1},\theta_{k})^{\prime}s$ are inactive, then we have $\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})<0$, hence, $f_{j}(x_{k+1},\theta_{k})<-\frac{\lambda_{k}^{(j)}}{\rho_{k}}.$ Thus, we have $\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}=-\lambda_{k}^{(j)}.$ Therefore, in inactive case also we have $\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}=\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}.$ Since the above equality holds for both active and inactive cases, we can write in full vector notation that $\lambda_{k+1}-\lambda_{k}=\rho_{k}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k}).$ Now by adding and subtracting $\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$, we get
@@
where in the last equality, we define $e_{k}=\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)-\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)$. For any $\lambda\in\mathbb{R}_{+}^{J}$, using the three-point identity and dividing both sides by $\rho_{k}$ we get
@@
$\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}$ $=\frac{1}{2\rho_{k}}\|\lambda_{k}-\lambda\|^{2}-\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda\|^{2}+(\lambda_{k+1}-\lambda)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+(\lambda_{k+1}-\lambda)^{\top}e_{k}$
@@
Since $a=b+c$, squaring both sides and using the fact that $(b+c)^{2}\geq\frac{1}{2}b^{2}-c^{2}$, we obtain
$\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}$ $=\rho_{k}^{2}\left(f_{j}(x_{k+1},\theta_{k})\right)^{2}$
$\geq\rho_{k}^{2}\left(\frac{1}{2}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-\left(f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})\right)^{2}\right).$
@@
$\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}\geq\rho_{k}^{2}\left[\frac{1}{2}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}\right].$
@@
$\frac{1}{2\rho_{k}}\sum_{j\in J_{k}^{+}}\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}\geq\frac{\rho_{k}}{4}\sum_{j\in J_{k}^{+}}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}.$
If $j\in J_{k}^{-}$, then $\lambda^{(j)}_{k+1}=0\implies\left(\lambda^{(j)}_{k+1}-\lambda^{(j)}_{k}\right)^{2}=\left(\lambda^{(j)}_{k}\right)^{2}$. Multiplying both sides by $\frac{1}{2\rho_{k}}$, and combining both of the above inequalities, we get
$\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}\geq\tfrac{\rho_{k}}{4}\sum_{j\in J_{k}^{+}}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}+\tfrac{1}{2\rho_{k}}\sum_{j\in J_{k}^{-}}(\lambda_{k}^{(j)})^{2}-\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}.$
@@
$-\left(\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}\right)\leq-\Biggl[\tfrac{\rho_{k}}{4}\sum_{j\in J_{k}^{+}}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}+\tfrac{1}{2\rho_{k}}\sum_{j\in J_{k}^{-}}(\lambda_{k}^{(j)})^{2}\Biggr].$ (11)
Now, notice that
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)$ $=\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)+\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)$
$=\sum_{j\in J_{k}^{+}}\Bigl(\frac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta_{k})\bigr)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta_{k})\Bigr)-\sum_{j\in J_{k}^{-}}\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}$
$\qquad+\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr).$
For each $j\in J_{k}^{+}$, we add and subtract the term $\tfrac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta^{*})$ inside the summation, hence we obtain
$\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$ $=\sum_{j\in J_{k}^{+}}\Bigl(\tfrac{\rho_{k}}{2}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta^{*})\Bigr)-\sum_{j\in J_{k}^{-}}\frac{(\lambda_{k}^{(j)})^{2}}{2\rho_{k}}+E_{k},$ (12)
where $E_{k}\triangleq\sum_{j\in J_{k}^{+}}\left(\tfrac{\rho_{k}}{2}\bigl(\left(f_{j}(x_{k+1},\theta_{k})\right)^{2}-\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}\bigr)+\lambda_{k}^{(j)}\bigl(f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})\bigr)\right)+\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)$. Now using the facts from (6) and (12), we can expand as follows
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
@@
$\qquad-\sum_{j\in J_{k}^{-}}\bigl(\lambda^{(j)}_{k+1}-\lambda^{(j)}\bigr)\Bigl(-\tfrac{\lambda^{(j)}_{k}}{\rho_{k}}\Bigr)-\sum_{j\in J_{k}^{+}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\sum_{j\in J_{k}^{-}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})+E_{k}.$
@@
$\quad-\sum_{j\in J_{k}^{+}}\left(\lambda^{(j)}_{k}+\rho_{k}f_{j}(x_{k+1},\theta^{*})-\lambda^{(j)}\right)f_{j}(x_{k+1},\theta^{*})-\sum_{j\in J_{k}^{-}}\left(-\lambda^{(j)}\right)\left(-\frac{\lambda^{(j)}_{k}}{\rho_{k}}\right)-\sum_{j\in J^{+}_{k}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})+E_{k}.$
@@
$\quad-\left(\lambda^{(j)}_{k}+\rho_{k}f_{j}(x_{k+1},\theta^{*})-\lambda^{(j)}\right)f_{j}(x_{k+1},\theta^{*})\Bigr]-\sum_{j\in J_{k}^{-}}\left(\tfrac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}+\lambda^{(j)}\Bigl(f_{j}(x_{k+1},\theta^{*})+\tfrac{\lambda_{k}^{(j)}}{\rho_{k}}\Bigr)\right)+E_{k}.$
@@
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad=-\sum_{j\in J_{k}^{+}}\tfrac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}-\sum_{j\in J_{k}^{-}}\left(\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}+\lambda^{(j)}\Bigl(f_{j}(x_{k+1},\theta^{*})+\frac{\lambda_{k}^{(j)}}{\rho_{k}}\Bigr)\right)+E_{k}.$ (13)